Theoretical study of avoided resonance overlapping and trapping effect in predissociation processes.

M. DESOUTER-LECOMTE, V. BREMS, Département de Chimie, Université de Liège, B6, Sart Tilman, B-4000 Liège, BELGIUM; \underline{J. Liévin}, Laboratoire de Chimie Physique Moléculaire, Université Libre de Bruxelles, CPI 160/09, B-1050 Bruxelles, BELGIUM.

The interaction of overlapping resonances and the resulting segregation between trapped and short-lived decay modes are studied in model cases of vibrational and electronic predissociations. Departure from the energy independent effective Hamiltonian formalism is shown to be decisive for a proper description of the migration of the resonance energies and widths when the strength of the external coupling is increased. Three different methods are used in a complementary way for analyzing this behaviour: the energy dependent partitioning method, the complex scaling method and the wavepacket propagation technique. The transition from the diabatic to the adiabatic regime is discussed in the electronic predissociation case.